OpenStudy (anonymous):

A regulation basketball has a diameter of 9.4 inches. What is the volume of the ball?

OpenStudy (anonymous):

V = (4/3)πr³ V = (4/3)π(4.5)³ V = (4/3)π(91.125) Using 3.1415 again: V = 381.69 So about 382 in³

OpenStudy (anonymous):

the radius in 4.7...

OpenStudy (anonymous):

then where the 4.5 is place 4.7 instead

OpenStudy (anonymous):

V = (4/3)πr³ V = (4/3)π(4.7)³ V = (4/3)π(103.823)

OpenStudy (anonymous):

i for some reason got 434.8